# Visual Whole-Body Control for Legged Loco-Manipulation

### v4

@@ Acknowledgments
We would like to thank Unitree for their technical support. We thank Chengzhe Jia for the great help on 3D modeling. We also thank Haichuan Che, Jun Wang, Shiqi Yang, and Jialong Li for their kind help in recording demos.
Appendix
I Notations and Background
I-A Reinforcement Learning
@@
I-B Proximal Policy Optimization
Proximal policy optimization (PPO) is one of the popular algorithms that solve RL problems. The basic idea behind PPO is to maximize a surrogate objective that constrains the size of the policy update. In particular, PPO optimizes the following objective:
@@ I-C Imitation Learning
I-D Whole-Body Control
Whole-body control in robotics refers to controlling a mobile manipulator with the function of locomotion and manipulation by one unified framework. In this work, we consider a system that contains a quadruped robot combined with a 6-DoF robot arm, each of which has an underlying PD controller, and we use a learned policy to generate control signals and feed them to the controller given 1) the robot states and 2) commands.
In detail, the policy used for controlling the quadruped robot is driven by external commands, such as the base velocity command; and the arm is controlled by a PD controller driven by the $\mathbb{S}\mathbb{E}(3)$ end-effector position-orientation command.
TABLE II: Details of low-level policy.
@@ I-D Whole-Body Control
Simulation We use the GPU-based Isaac Gym simulator [47] where we train 4096 environments in parallel.
II Training Details
II-A Low-Level Training
Low-level commands sampling. To train a robust low-level controller for goal-reaching, we should sample and track various goals (i.e., base velocity commands and end-effector commands) as possible. As for the base velocity commands, we uniformly sample the linear velocity from a range of $[-0.6,0.6]$, and the angular velocity from a range of $[1.0,1.0]$. Regarding the end-effector, we employ a height-invariant coordinate system originating at the robot’s base. Such a design ensures the sampled trajectories are always on a sphere, guaranteeing the consistency of these samples; besides, it prevents the sampled goals from being affected by the quadruped robot’s height. In particular, we interpolate the end-effector goal command $\mathbf{p}^{\text{cmd}}$ between the current end-effector position $\mathbf{p}$ and a randomly selected end-effector position $\mathbf{p}^{\text{end}}$, at intervals of $T_{\text{traj}}$ seconds:
@@ II-A Low-Level Training
$\mathbf{p}^{\text{end}}$ is resampled in a fixed timestep or if any $\mathbf{p}_{t}^{\text{cmd}}$ leads to self-collision or collision with the ground, similar to the linear and yaw velocity commands $v_{\text{lin}}^{\text{cmd}}$, $\omega_{yaw}^{\text{cmd}}$ that are uniformly sampled from a given range. Fig. 13 shows a screenshot of the low-level training, where the robot is required to track the given velocity commands and the end-effector pose.
Figure 13: A screenshot of low-level training, where the robot is required to track the given velocity commands (linear and yaw) and the end-effector pose.
@@
Regularized online adaptation for sim-to-real transfer. We take the training paradigm of Regularized Online Adaptation (ROA) [22, 12, 32] in the RL training to help decrease the realizability gap when deploying our system into the real world. As illustrated in Fig. 3, we have an encoder $\mu$ that takes the privileged information $e$ such as friction and mass as input and predicts an environment extrinsic latent vector $z^{\mu}$. Then, the adaptation module $\phi$ estimates this $z^{\mu}$ only based on recent history low-level observations. The adaptation module $\phi$ is trained by imitating $z^{\mu}$ online. We can then formulate the loss function of our whole low-level policy w.r.t. policy’s parameters $\theta_{\pi^{\text{low}}}$, privileged information encoder’s parameters $\theta_{\mu}$, and adaptation module’s parameters $\theta_{\phi}$ as:
@@
$\begin{split}L\left(\theta_{\pi^{\text{low}}},\theta_{\mu},\theta_{\phi}\right)=-L^{PPO}(\theta_{\pi^{\text{low}}},\theta_{\mu})&+\lambda\left\|z^{\mu}-\operatorname{sg}\left[z^{\phi}\right]\right\|_{2}\\ &+\left\|\operatorname{sg}\left[z^{\mu}\right]-z^{\phi}\right\|_{2}\end{split}$ (7)
@@
where $\operatorname{sg}\left[\cdot\right]$ is the gradient stop operator, and $\lambda$ is the Laguagrian multiplier. We optimize this loss function using the dual gradient descent method, and [22] has identified that such an optimization converges under mild conditions. It is worth noticing that RMA [36] is a special case of ROA, where the Laguagrian multiplier $\lambda$ is set to be constant zero. The adaptation module $\phi$ starts training only after convergence of the policy $\pi^{\text{low}}$ and the encoder $\mu$.
Reward functions. The detailed definitions of the low-level reward functions are shown in Tab. II, where we define $\phi=\exp(-\frac{\|x\|^{2}}{0.25})$ for tracking the desired velocity of the robot. It is worth noting that we introduce the swing phase tracking rewards to help the robot learn trotting, and $C_{\text{foot}}^{\text{cmd}}(\theta^{\text{cmd}},t)$ computes the desired contact state of each foot from the phase and timing variable, as described in [48].
II-B High-Level Training
The privileged teacher policy is trained through RL, specifically, we use PPO for training with designed reward functions. Reward functions We put all reward functions along with their weights of training the high-level policies in Tab. III. Except for $r_{\text{progress}}$, $r_{\text{completion}}$ and $r_{\text{completion}}$ that is determined by the task stage, we define a set of assistant rewards to help the algorithm converge faster. Among them, $r_{\text{acc}}$ limits the arm joints velocities’ change rate; $r_{\text{cmd}}$ is designed to encourage the robot to slow its velocity when it is close to the object to be picked., $r_{\text{action}}$ helps smooth the output of the high-level policy to prevent it from changing quickly; $r_{\text{ee orn}}$ and $r_{\text{base orn}}$ encourages the arm and the robot body to be aimed at the object; $r_{\text{base approach}}$ encourages the robot to be close to the object when the robot is far away from the object.
@@ II-B High-Level Training
TABLE III: Details of High-Level Rewards, where $\dot{\mathbf{q}}$ denotes arm joint velocities; $v_{x}^{*}$ denotes base velocity command; $\mathbf{d}_{\text{obj}}$, $\mathbf{d}_{\text{ee}}$ and $\mathbf{d}_{\text{base}}$ are the direction from robot body to the object, the direction from end-effector to the object and the direction of the robot body orientation, separately; $\mathbf{x}_{\text{obj}}$ and $\mathbf{x}_{\text{base}}$ are the position coordinates in world frame of the object and the robot body.
@@
Stage Rewards $r_{\text{approach}}$ (Eq. (2)) $0.5$
$r_{\text{progress}}$ (Eq. (3)) $1.0$
$r_{\text{completion}}$ (Eq. (4)) $3.5$
Assistant Rewards $r_{\text{acc}}=1-\exp(\|\dot{\mathbf{q}}_{t-1}-\dot{\mathbf{q}}_{t}\|)$ $-0.001$
@@
$r_{\text{ee orn}}=|\mathbf{d}_{\text{obj}}\cdot\mathbf{d}_{\text{ee}}|$ $-0.05$
@@
$r_{\text{base approach}}=|\mathbf{x}_{\text{obj}}-\mathbf{x}_{\text{base}}|$ $0.01$
@@
Additional Training Techniques We design several training techniques with our prior knowledge to help the high-level policy to be well-performed and stable, not only in the simulation, but can also be transferred into the real world. We briefly introduce them in the following:
@@
    Command clip curriculum. We gradually clip the linear velocity command during training, as we find that a large velocity helps accelerate the learning convergence but a slow velocity highly stabilizes the real world performance.
@@
    Randomly changes the object position. During training, we take a small probability ($10\%$) to randomly change the object position and pose. This helps the robot to learn the retrying behaviors when it fails during one trial.

### v5

@@ Acknowledgments
This project was supported, in part, by the Amazon Research Award, the Intel Rising Star Faculty Award, and gifts from Qualcomm, Covariant, and Meta. We thank Chengzhe Jia for the great help on 3D modeling. We also thank Haichuan Che, Jun Wang, Shiqi Yang, and Jialong Li for their kind help in recording demos. We thank Xuanbin Peng for helping with the experiments.
Appendix A Extended Related Works
Appendix B Notations and Background
B.1 Reinforcement Learning
@@
B.2 Proximal Policy Optimization
Proximal policy optimization (PPO) [81] is one of the popular algorithms that solve RL problems. The basic idea behind PPO is to maximize a surrogate objective that constrains the size of the policy update. In particular, PPO optimizes the following objective:
@@ B.3 Imitation Learning
B.4 Whole-Body Control
Whole-body control in robotics refers to controlling a mobile manipulator with the function of locomotion and manipulation by one unified framework [9, 8]. In this work, we consider a system that contains a quadruped robot combined with a 6-DoF robot arm, each of which has an underlying PD controller, and we use a learned policy to generate control signals and feed them to the controller given 1) the robot states and 2) commands.
In detail, the policy used for controlling the quadruped robot is driven by external commands, such as the base velocity command; and the arm is controlled by a PD controller driven by the $\mathbb{S}\mathbb{E}(3)$ end-effector position-orientation command and inverse kinematics.
Table 2: Details of low-level policy.
@@ B.4 Whole-Body Control
Appendix C More Training Details
C.1 Low-Level Training
Reward functions. Our reward for training the low-level policy for whole-body control mainly consists of four parts: command-following reward, energy reward, alive reward, and phase reward. The command following reward encourages the robot to track and follow the provided commands and explore various whole-body behaviors; the energy reward penalizes energy consumption to enable smooth and reasonable motion; the alive reward encourages the robot not to fail; the phase reward encourages the robot to walk in a steady behavior as mentioned above. The detailed definitions of the low-level reward functions are shown in Tab. 2, where we define $\phi=\exp(-\frac{\|x\|^{2}}{0.25})$ for tracking the desired velocity of the robot. It is worth noting that we introduce the swing phase tracking rewards to help the robot learn trotting, and $C_{\text{foot}}^{\text{cmd}}(\theta^{\text{cmd}},t)$ computes the desired contact state of each foot from the phase and timing variable, as described in Margolis and Agrawal [59].
Low-level commands sampling. To train a robust low-level controller for goal-reaching, we should sample and track various goals (i.e., base velocity commands and end-effector commands) as possible. As for the base velocity commands, we uniformly sample the linear velocity from a range of $[-0.6m/s,0.6m/s]$, and the angular velocity from a range of $[1.0rad/s,1.0rad/s]$.
Regarding the end-effector, same to Fu et al. [8], we employ a height-roll-pitch-invariant coordinate system originating at the robot’s base, where the sampled goal poses are only affected by the robot’s horizontal movement instead of changes in height. In detail, the origin (cyan point in Fig. 9) is set on a fixed plane, which is on the same line of Z axis as the arm base. Therefore, the sampled goals are also not affected by the robot’s height, and orientation of the roll and pitch axis, but only by the orientation in the yaw axis. Such a design ensures the sampled trajectories are always on one local sphere at any time, guaranteeing the consistency of these samples. Besides, it helps to learn a whole-body behavior to control the legs and the body height to enable the arm to reach an expanded space of goals. If not, the robot may fail to reach a goal out of the workspace of the arm without bending or rising. Therefore, the body and the arm are not decoupled, although we use IK to control the arm.
We parameterize the end-effector position command $\mathbf{p}^{\text{cmd}}$ in spherical coordinate $(l,p,y)$, where $l,p,y$ lie in ranges of [0.4, 0.95], [-1 * $\pi$ / 2.5, 1 * $\pi$ / 3], [-1.2, 1.2], respectively. To make a smooth end-effector movement, we set $\mathbf{p}^{\text{cmd}}$ by interpolation between the current end-effector position $\mathbf{p}$ and a randomly sampled end-effector position $\mathbf{p}^{\text{end}}$, at intervals of $T_{\text{traj}}$ seconds:
@@ C.1 Low-Level Training
$\mathbf{p}^{\text{end}}$ is resampled in a fixed timestep or if any $\mathbf{p}_{t}^{\text{cmd}}$ leads to self-collision or collision with the ground. Fig. 9 shows screenshots of the low-level training, where the robot is required to track the given velocity commands and the end-effector pose. The orientation command $\mathbf{o}^{\text{cmd}}$ is uniformly sampled from $\mathbb{SO}(3)$ space.
Figure 9: Screenshots of low-level training, where the robot is required to track the given velocity commands (linear and yaw) and the end-effector pose. We employ a height-invariant coordinate system originating at the robot’s base, such that the sampled goal poses are only affected by the robot’s horizontal movement (translation in X, Y axes and orientation in the yaw axis). The cyan point represents the origin of the resampled goals; the blue point and the yellow point are the current end-effector position and the goal position, respectively. The red dots are the interpolated goals. The red lines represent the space for determining self-collision when resampling new goals.
@@
Regularized online adaptation for sim-to-real transfer. We take the training paradigm of Regularized Online Adaptation (ROA) [8, 84, 17] in the RL training, to help decrease the realizability gap when deploying our system into the real world. As illustrated in Fig. 10, in the first training phase, we train an encoder $\mu$ that takes the privileged information $e$ (such as friction and mass…etc.) as inputs, and predicts an environment extrinsic latent vector $z^{\mu}$, which will be used for policy inference. Note that the privileged information is not available in real robot deployment. Thus we will need another way to estimate the latent vector $z^{\mu}$. To this end, we train the adaptation module $\phi$ in a second training phase, to imitate/distill the output of the encoder $\mu$, and estimate this $z^{\mu}$ only based on recent history low-level observations. The loss function of the low-level policy w.r.t. policy’s parameters $\theta_{\pi^{\text{low}}}$, privileged information encoder’s parameters $\theta_{\mu}$, and adaptation module’s parameters $\theta_{\phi}$, can be then formulated as:
@@
$\begin{split}L\left(\theta_{\pi^{\text{low}}},\theta_{\mu},\theta_{\phi}\right)=-L^{PPO}(\theta_{\pi^{\text{low}}},\theta_{\mu})&+\lambda\left\|z^{\mu}-\operatorname{sg}\left[z^{\phi}\right]\right\|_{2}+\left\|\operatorname{sg}\left[z^{\mu}\right]-z^{\phi}\right\|_{2}\end{split}$ (3)
@@
where $\operatorname{sg}\left[\cdot\right]$ is the gradient stop operator, and $\lambda$ is the Laguagrian multiplier. It is worth noticing that RMA [52] is a special case of ROA, where the Laguagrian multiplier $\lambda$ is set to be constant zero. The adaptation module $\phi$ starts training only after convergence of the policy $\pi^{\text{low}}$ and the encoder $\mu$.
Figure 10: Regularized Online Adaptation (ROA) training of the low-level policy.
Domain randomization. When training, we randomize the terrains, including flat planes and rough ones. We also randomized the friction between robots and terrains. Besides, we randomized the quadruped’s mass and center of mass. These randomizations, together with our ROA module, contribute to a robust low-level policy.
C.2 High-Level Training
Reward functions. The task-specific reward function for training our state-based policy comprises stage rewards and assistant rewards. 1) We design three stages for a task. The first stage is approaching, and the corresponding reward $r_{\text{approach}}$ encourages the gripper and quadruped to get close to the target object: $r_{\text{approach}}=\min(d_{\text{closest}}-d,0)~{},$where $d_{\text{closest}}$ is the current minimum distance between the gripper and the target object. The second stage is task progress, and the reward $r_{\text{progress}}$ leads the robot to the desired final state, e.g., lifting the object in the pickup task: $r_{\text{progress}}^{\text{pickup}}=\min(d-d_{\text{highest}},0)~{},$where $d_{\text{highest}}$ is the current largest height of the object. The third stage is denoted task completion, and reward $r_{\text{completion}}$ bonus when the agent completes the task. $r_{\text{completion}}=\mathbf{1}(\text{Task is completed}).$For example, in the pickup task, the task completion condition is that the object height exceeds the required threshold. Note that at one time, the robot can only be at one stage and receive one stage reward of the above three, ensuring that different task stages are not intervened in. The assistant rewards are designed to a) smooth the robot’s behavior to be easier to transfer to the real world, like penalty huge action commands and joint acceleration; b) prevent deviation and sampling useless data, such as acquiring the robot’s body and gripper to point to the object, which guarantees that the camera does not lose tracking objects. 2) Beyond these stage rewards, we also design a set of assistant reward functions that help accelerate the learning procedure. We put all reward functions along with their weights of training the high-level policies in Tab. 3. Except for $r_{\text{progress}}$, $r_{\text{completion}}$ and $r_{\text{completion}}$ that is determined by the task stage, we define a set of assistant rewards to help the algorithm converge faster. Among them, $r_{\text{acc}}$ limits the arm joints velocities’ change rate; $r_{\text{cmd}}$ is designed to encourage the robot to slow its velocity when it is close to the object to be picked., $r_{\text{action}}$ helps smooth the output of the high-level policy to prevent it from changing quickly; $r_{\text{ee orn}}$ and $r_{\text{base orn}}$ encourages the arm and the robot body to be aimed at the object; $r_{\text{base approach}}$ encourages the robot to be close to the object when the robot is far away from the object.
@@ C.2 High-Level Training
Table 3: Details of high-level rewards, where $\dot{\mathbf{q}}$ denotes arm joint velocities; $v_{x}^{*}$ denotes base velocity command; $\mathbf{d}_{\text{obj}}$, $\mathbf{d}_{\text{ee}}$ and $\mathbf{d}_{\text{base}}$ are the direction from robot body to the object, the direction from end-effector to the object and the direction of the robot body orientation, separately; $\mathbf{x}_{\text{obj}}$ and $\mathbf{x}_{\text{base}}$ are the position coordinates in world frame of the object and the robot body. $r_{\text{approach}}$, $r_{\text{progress}}$, and $r_{\text{completion}}$ are defined at Section 3.2.
@@
Stage Rewards $r_{\text{approach}}$ $0.5$
$r_{\text{progress}}$ $1.0$
$r_{\text{completion}}$ $3.5$
Assistant Rewards $r_{\text{acc}}=1-\exp(-\|\dot{\mathbf{q}}_{t-1}-\dot{\mathbf{q}}_{t}\|)$ $-0.001$
@@
$r_{\text{ee orn}}=\text{cos}(\mathbf{d}_{\text{obj}}\cdot\mathbf{d}_{\text{ee}})$ $0.01$
@@
$r_{\text{base approach}}=(1+\text{tanh}(-10*\|x_{obj}–x_{base}-0.6\|))$ $0.01$
Teacher policy network architecture. Our privileged teacher high-level policy network includes a pre-trained PointNet encoder for modeling object pointclouds, followed by a two-layer MLP that gives the high-level actions.
Student policy network architecture. Our vision-based high-level policy network comprises a CNN and a two-layer MLP that outputs the high-level actions. We concatenate the images and other states as observations. We stack 4-step images which are passed through the two-layer CNN network with kernel size being $5$ and $3$, respectively, and then encoded as a $64$-dim latent vector. Then, we concatenate this latent vector with other states as a flattened vector. This vector is then passed to a two-layer MLP with a hidden size of 64. We use ELU as the activation.
@@
Additional training techniques. We design several training techniques with our prior knowledge to help the high-level policy to be well-performed and stable, not only in the simulation, but can also be transferred into the real world. We briefly introduce them in the following:
@@
    Command clip curriculum. We gradually clip the linear velocity command during training, as we find that a large velocity helps accelerate the learning convergence but a slow velocity ensures a safe, stable, and smooth behavior in the real world.
@@
    Randomly changing the object position. During training, we take a small probability ($10\%$) to randomly change the object position and pose. This helps the robot to learn the retrying behaviors when it fails during one trial.
@@
Appendix D Summary of Domain Randomization
We summarize and list all domain randomization here during the simulated training. Low-level policy. The randomization of the low-level policy is to train a robot with robust walking skills that adapt to different surroundings, while also knowing how to adjust body poses to reach different end-effector goals.
  • Terrains, including flat planes and rough ones.
  • Friction between robots and terrains.
  • Mass of the quadruped and the arm.
  • Center of mass of the robot.
  • Motor strength.
High-level teacher policy. The randomization of the high-level part mainly focuses on learning the manipulation task with generalization. The high-level teacher policy follows the same randomization as the low-level policy, along with the following additional terms:
  • High-level frequency by involving random low-level policy calls during one high-level call. In particular, we randomize 6-8 calls considering the high-level inference time and signal transmission latency, corresponding to 0.12-0.16s (low-level policy run at 50 Hz), i.e., 6.25-8.33Hz.
  • Kp and Kd parameters for the arm controller.
  • Table heights, 0-0.5m.
  • Initial position and pose of the robot.
  • Position and pose of objects.
High-level student policy. The high-level student policy follows the same randomization as the teacher policy, along with the following additional term:
  • Camera latency.
Appendix E Motivations for IK-based Arm Control
In this work, we control the arm by solving the target joints with inverse kinematics, given the predicted target end-effector 6-DoF poses, rather than directly predicting the arm joints. Such a design is motivated by lessons learned from the previous work, DeepWBC [8], which struggles to control the accurate end-effector position and orientation. To verify the advantage of our IK-based control predicting end-effector poses compared to predicting the target joints, we conduct a goal-reaching experiment, in which we randomly sample end-effector goals as in training, and end the episode once the goal is reached (the distance of end-effector and the goal is less than 2cm) or the episode time is out. We quantitatively evaluate the average error of position and orientation in Tab. 4. The results demonstrated a clearly better precision compared to the target joint prediction solution in Fu et al. [8], which contributes a lot to training an effective high-level policy to complete complicated manipulation tasks.
Approach Status Position Error (m) $\downarrow$ Orientation Error (rad) $\downarrow$
Mean Min Max Mean Min Max
Joint prediction ([8]) Standing $0.236\pm 0.0$ $0.146\pm 0.0$ $0.325\pm 0.0$ $1.154\pm 0.0$ $0.616\pm 0.0$ $1.686\pm 0.0$
EE prediction+IK (ours) $\mathbf{0.047\pm 0.0}$ $\mathbf{0.027\pm 0.0}$ $\mathbf{0.067\pm 0.0}$ $\mathbf{0.850\pm 0.01}$ $\mathbf{0.160\pm 0.0}$ $\mathbf{2.269\pm 0.05}$
Joint prediction ([8]) Walking $0.414\pm 0.0$ $0.182\pm 0.0$ $0.558\pm 0.0$ $2.015\pm 0.0$ $1.165\pm 0.03$ $3.198\pm 0.04$
EE prediction+IK (ours) $\mathbf{0.029\pm 0.0}$ $\mathbf{0.008\pm 0.0}$ $\mathbf{0.050\pm 0.0}$ $\mathbf{0.332\pm 0.02}$ $\mathbf{0.071\pm 0.03}$ $\mathbf{0.743\pm 0.08}$
Table 4: Precision of reaching goals, evaluated in simulation with 3 seeds, each over 100 episodes. We record the results while the robots are standing and walking. In both status our solution achieves the best performance.
Appendix F Sim-Real Visual Observation Comparison
We compare the trajectories and segmented depth images in simulation and real-world sampled by VBC, shown in Fig. 11.
Figure 11: Comparison between trajectories and segmented depth images in simulation and real-world sampled by VBC. Left: The real-world behaviors match well with the simulation, indicating a less sim-to-real gap in our method. Right: Our TrackingSAM-based working pipeline allows us to obtain the segmentation mask precisely during deployment (not the same object observed). The depth images are pre-processed by hole-filling and clipping.
Appendix G Real World System
Hardware setup. As visualized in Fig. 4, we mount two cameras on our robot platform, which are both RealSense D435 cameras. B1 itself has an onboard computer, and we also attached an additional onboard computer, Nvidia Jetson Orin, to the robot. The B1, Z1, and the Orin computer are all powered by B1’s onboard battery. The inference of the low-level policy and the control of the arm is conducted at the onboard computer of B1, and that of the high-level vision-based planning policy is made on Nvidia Jetson Orin. It is worth noting that the computation and the power of our robot system are fully onboard and, therefore, untethered by any PCs.
Software system design. During deployment, the low-level policy that controls the quadruped works at a frequency of 50Hz; in the meantime, the high-level policy runs at $\sim$8Hz, which is more than five times slower than the low-level policy; besides, the arm is required to work at a frequency of more than 800Hz. Since these policies work on different computers, along with the quadruped and the arm should work under different frequencies, we design a client-server multi-processing system for this work, as shown in Fig. 12. We set up three processes at the low-level policy side: a quadruped robot control process, an arm control process, and a goal receiver process behaving as a client. The high-level policy combines the vision inputs from the depth cameras and the information from the low-level request as the observation works as a server. Every 0.1 seconds, the goal receiver sends a request with the necessary information to the server; the high-level policy makes the latest inference and sends the high-level action back to the robot; at the same time, the quadruped and the arm control processes gather the latest high-level commands to execute, while keep updating the necessary information for the goal receiver process.
Component runtime analysis. We record the runtime of all our components running in real-world experiments, shown in Tab. 5. Before an autonomous working pipeline, users are required to click on images to annotate the object to be grasped with SAM, which is further processed by AOT to initialize the first reference frame to track. After initialization, the high-level process reads in a new frame from two cameras, and pre-process them by resizing, hole-filling, and clipping; the images will be further processed by TrackingSAM to track and provide segmentation masks; the masks and segmented images are concatenated as observations along with the proprioception states sent by the low-level process, and call the high-level model to predict the command. The high-level process finally sends the command back to the low-level process, which calls the low-level model to predict the leg joints. Note that we did not record the time cost of process communication between the high-level and the low-level, each of which runs on different devices and communicates through HTTP. During training, we gave a slightly larger randomization range than the recorded total time (also see in Appendix D). Overall, the high-level runtime limits the control frequency of the high-level policy to 6-8Hz. We also fix the low-level control frequency as 50Hz, even if the policy inference is rather fast.
High-Level Low-Level
TrackingSAM Pre-Process Model Inference Model Inference
SAM clicking AOT Init AOT Tracking
Time (s) $1.369\pm 0.33$ $0.323\pm 0.72$ $0.083\pm 0.0$ $0.031\pm 0.01$ $0.003\pm 0.0$ $0.011\pm 0.0$
Stage Initialization Autonumous
Device External NVIDIA Jetson Orin 64GB Internal NVIDIA Jetson Xavier NX
Table 5: Component runtime analysis. All components are run on real robots, and the results are averaged over 100 times.
Appendix H Details of Used Objects
Figure 13: Examples of simulated training objects and real-world objects.
For the simulation, we included $34$ objects adapted from the YCB dataset [28]. According to the object shapes, we roughly divided them into $7$ categories: ball, long box, square box, bottle, cup, bowl and drill. In the real-world experiments, we choose 14 objects, including four irregular objects, six common objects, and four regular shapes. Objects are illustrated in Fig. 13
Appendix I Failure Analysis and Future Works
Although we showed great success in picking various objects, we found several failure cases due to the limitations of the system: 1) Compounding error. Our pipeline is hierarchical and requires each module to be coherent and precise to make the robot work well. 2) In-precise depth estimation. The RGBD camera does not reflect a perfect depth estimation, especially on reflective objects. 3) Bad gripper. The gripper of the Unitree Z1 Arm is a beak-like gripper instead of a parallel one, which tends to push objects away and is hard for accurate manipulation. 4) Unstable vision inputs. The tracking SAM that is used to provide the mask of annotated objects, may lose track due to camera deviation and occlusion, or be confused when the selected objects have a similar color to the other things. This is the most common cause leading to failure in our real-world experiments. In our current pipeline, before the robot autonomously picks up objects, we must manually specify the target object in both cameras mounted on the robot. Thanks to the generalizability of trackingSAM, we find that as long as the object is visible to one camera (either the head or the wrist camera), the annotation mask can be propagated to another view for automated picking, and free the robot from being specified in the between of the range of the two cameras, which improves the usability of our system. Since the requirements of annotation and object visibility are totally decoupled from our sim-to-real framework, in the future we can further improve and replace these modules with automated annotation algorithms and a search/exploration module to further improve the automation of the whole system. Our choice of visual inputs—specifically masks and segmented depth images—deliberately excludes RGB images that provide color information. This limitation restricts the system from performing tasks that rely on color differentiation, such as distinguishing between Fanta and Sprite cans, sorting garbage by category, or pressing a green button. During the training of our high-level policy, we utilized 33 distinct objects. While this approach demonstrates strong generalization capabilities on unseen objects in real-world scenarios, future work could enhance this by introducing randomized object sizes to further improve generalization. Future works should also consider upgrading the hardware, like replacing a better gripper and RGBD cameras that provide better depth estimation; and simplify the working pipeline.
Appendix J Additional Details of TrackingSAM
TrackingSAM [24] provides a very simple and intuitive way to provide segmentation masks which only requires minimal manual intervention. It invokes SAM [26] to generate an initial segmentation mask of objects before the tracking begins, where the user only needs to make a few clicks (as shown in Fig. 14). Given the initial segmentation, we use it to populate two instances of trackers for images from the wrist camera and the head camera, respectively. The tracker used is AOT (Associating Objects with Transformers) [25], using the most lightweight AOT-Tiny variant, which is implemented based on MobileNet and can achieve nearly 10 fps on the onboard computer.
Figure 14: Segmented examples of the TrackingSAM tool.